OpenStudy (anonymous):

If all the dimensions of a cube are doubled, then the new volume is ____ times greater.

ganeshie8 (ganeshie8):

volume = s^3 if you double the side, new volume = (2s)^3 = 2^3 x s^3

ganeshie8 (ganeshie8):

figured out how many times ?

OpenStudy (anonymous):

No

OpenStudy (whpalmer4):

Think of plugging the scale factor for each dimension into the formula. The formula for the volume of a cube is V = l*w*h, right? Well, if you double each of the dimensions, your new V = 2*2*2 * the old V, or 8x. Similarly, say you have a rectangle, and you make it 3x as long and 2x as wide. Area of the rectangle is A = l*w, so the new rectangle is 3*2* the old A, or 6x. Make sense?

OpenStudy (whpalmer4):

The example I personally find most valuable is for the area of the circle. The area of the circle is \(A = \pi r^2\). If you double the radius of the pizza, the new pizza has (2)^2 = 4 the area of the old pizza, and that way there ought to be enough for me :-)